Shift summary for August 22, 1999 DAY shift

Vivian O'Dell (odell@fnal.gov)
Sun, 22 Aug 1999 15:34:00 -0500

Shift summary for DAY shift, 8/22/99

On Shift: T. Yamanaka, V. O'Dell

Hours of beam available at all intensities: 8.5h
Hours of beam available at nominal intensity: 8.5h
Hours of E832 data (special runs not included): 8.0h (rounded off)
Hours of special runs (laser scans, muon runs...):0

Accounting of lost time: tape swap, kfrends
========================================================================
Runs ended this shift:

14299 - 34.4 K vac 2pi0

Tapes brought to FCC this shift: QK(S,C,K)306-308 QKB107

========================================================================
Accesses during shift: None. Managed to keep Taku out of the hall.
========================================================================
Overview of shift: smooth running at ~6.8 E12 with sporadic, intensive
paging. Common theme of "Don't worry, be happy."
========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
14299 141 07:00 Day shift turns up. One member scoffs down lots
of chocolate chip cookies left by last shift.

14299 08:00 Shift member on sugar high begins a constant stream
of paging:

1) paged Hogan about poor track link efficiency
plot for beam region. Hogan said:
"We know there is a problem, but there isn't
much we can do about it."

2) paged JasonL about banana spottiness above
normal in banana summary plots. Jason said:
"Don't worry until it goes above the warning
mark."

3) paged RickK about crappy looking cosmic
counter illuminations in the cosmic#summary
plots. Rick said he thinks the pedestals are
wrong and not to worry. When further pressed
about possible dead channels, he said
"I don't care."

4) paged Elizabeth about dead channel in BA
(channel 106 in fpedcal#summary plots). She
said "Don't worry about it."

5) paged Doug about vacuum pump oil, because
both people on shift are too old to see the
pump oil level on the camera. (JasonL should
make sure old people are teamed up with young
sharp-eyed people.) Doug will probably
say "Don't worry about it", but we can only
speculate at this point.

14299 08:35 Ed calls and says he has adjusted the gain
for the CsI channel with a large PMT gain
drift wrt database (large meaning 2.5%).
When one shift member presses the other on
which CsI channel this is, the other points
to the online plot on the CsI histograms screen
showing an outlier in the PMT gain drift and
says "That one". When further pressed, he gets
testy and says "Ed knows which one it is." At
any rate, we shouldn't worry about it and the
plot should look fine when we start the next
run.

14299 233-237 08:53 ktevana errors, busy dyc 53. Taku reseats
cables for UCLA latch and fixes the problem.
We lose about 4 spills.

14299 09:15 Look up above offensive CsI channel. It is
channel 1382.

14299 09:30 Effects of sugar start wearing off. It's
going to be a loooong day.

14299 09:50 Taku makes crank pager calls. Who is
722-6998?

14299 306 10:25 The quiet monotony of beaming palm pilot
applications is shattered by Jim Grahm's voice
demanding a tape mount. Toyed with idea
of paging him, but just mounted the tape
instead.

14299 343-345 11:15 DYC branch 4 crate 23 busy. A couple of
kfrends fixes it.

14299 11:30 Doug returns page and says
"Don't worry about it."

14299 360-365 11:40 ktevana errors cured by kfrending.
Why does ktevana complain about missing
F words anyway? Does ktevana expect crates
to use lots of F words? Also, kfrend seemed
to take an inordinately long time (perhaps it
collided with other processes and used the
F word once too often...). I certainly would
not recommend using any F words on the
Eisenhower.

14299 last 12:23 "Boy, I'm full." Swap tapes. Crack team
does it in record time.

14300 1 12:33 start of new run. Taku forces innocent
student to fill in paper work, leaving
shift crew free to run amuck.

14300 13:00 Rogue channel 1382 is no longer rogue,
as promised by Ed.

14300 13:30 Dave Northaker fills alcohol bubbler.

14300 14:15 Taku & student leave to check gasses.

14300 14:40 Taku & student still gone. Viv afraid to
check on them in case of a repeat of the
Blair Witch project...

14300 14:45 Taku & student return unscathed.

14300 121 15:15 Laser crate hangs readout. Taku on the
job.

14300 123 15:17 Taku's magic works and we're live again.

14300 15:20 Taku insists on paging Sydney, and Viv is
powerless to stop him.

14300 15:25 Sydney calls in. Taku is impressed by his
response time and explains to Sydney that
we're just bored. Sydney is not amused.
Screams of a child are heard in the background
and when Sydney is quizzed on the identity
of the child, he responds (exasperated):
"I really couldn't tell you."

14300 15:25 JasonL arrives.

14300 15:32 Where is John Shields? Day crew gives up
and leaves.

========================================================================
Follow up Notes on Previous Shift
========================================================================
Items to Follow up on Next Shift
========================================================================